### v1

# DPBridge: Latent Diffusion Bridge for Dense Prediction

Abstract
Diffusion models have demonstrated remarkable success in dense prediction problems, which aims to model per-pixel relationship between RGB images and dense signal maps, thanks to their ability to effectively capture complex data distributions. However, initiating the reverse sampling trajectory from uninformative noise prior introduces limitations such as degraded performance and slow inference speed. In this work, we propose DPBridge, a generative framework that formulates dense prediction tasks as image-conditioned generation problems and establishes a direct mapping between input image and its corresponding dense map based on fully-tractable diffusion bridge process. This approach addresses aforementioned limitations in conventional diffusion-based solutions. In addition, we introduce finetuning strategies to adapt our model from pretrained image diffusion backbone, leveraging its rich visual prior knowledge to facilitate both efficient training and robust generalization ability. Experimental results shows that our DPBridge can achieve competitive performance compared to both feed-forward and diffusion-based approaches across various benchmarks, highlighting its effectiveness and adaptability.
Figure 1: We present DPBridge to tackle general dense prediction tasks by establishing a stochastic transport between an RGB image and its corresponding dense map in the latent space based on a tractable diffusion bridge process. In addition, we introduce finetuning strategies to adapt large pretrained diffusion backbone to construct our bridge, leveraging its rich prior knowledge to enable both efficient training and robust generalization.
1 Introduction
Diffusion models recently emerge as a family of powerful generative models that are able to capture complex data distributions through a sequence of forward and reverse processes [26, 46]. They accelerate the development of modern generative systems, and have consistently obtained state-of-the-art performance across various domains, such as image synthesis and video generation [17, 46, 27], controlled editing [61, 5, 50], human pose estimation [29, 30, 28], inverse problem solving [11, 53, 40].
Despite such success, diffusion models still face limitations that hinder broader applications, especially in discriminative tasks like dense prediction. One major limitation stems from the fact that standard diffusion reverse process always starts from Gaussian noise, which contains little information about the target distribution. This will lead to degraded performance or prolonged convergence times. Current diffusion-based approaches formulate this type of tasks as image-conditioned generation problems and adopt measures like incorporating external conditions to the model [47, 61, 25], modifying the generation pipeline [41, 54], or adding extra guidance terms during sampling [11, 53] to deal with this issue. In this way they hope to better model the conditional generation process.
Instead of starting from noise, an feasible alternative is to leverage the inherent problem structure. Since the goal of dense prediction tasks is to model the probability transport between image and dense signal map distributions, it’s a rather straightforward idea to directly start the generative process from input RGB images, as they are structurally correlated to target signal maps and can naturally serve as more informative prior compared to noise. This motivates the exploration of a new family of computational frameworks, i.e. diffusion bridge models [39, 63, 36, 37]. They are based on a special type of diffusion process conditioned on given endpoints, thus enabling optimal transport between two distributions and generalizing beyond Gaussian priors.
Inspired by these intuition, we propose Dense Prediction Bridge (DPBridge), a latent diffusion bridge model adapted from pretrained image diffusion backbone to handle general dense prediction tasks though the image-conditioned generation pipeline. In essence, we construct diffusion bridge process between the latent distributions of images and dense signal maps, with all intermediate samples along the trajectory being convex combinations of the paired data. The generative process thus becomes a series of transformations that iteratively synthesize desired signal maps starting from input images rather than pure noise, enabling faster inference speed as well as better generation quality.
In addition, to better adapt pretrained diffusion backbone to our DPBridge construction, we also propose novel finetuning strategies to allow for more effective knowledge transfer. In this way our model is augmented with rich visual prior knowledge, which not only helps boosting the perception performance as well as generalization capability, but also significantly speeds up training convergence. Fig.(1) is a high-level overview of DPBridge, while also demonstrating its effectiveness on three representative dense prediction tasks, i.e. depth estimation, surface normal prediction, and semantic segmentation, proving that our framework is theoretically universal and can be instantiated to different scenarios.
To summarize, our contributions are as follows:
  • We propose DPBridge framework to deal with dense prediction tasks with generative pipeline by formulating them as instances of latent diffusion bridges.
  • We propose finetuning strategies to effectively adapt large pretrained backbone to diffusion bridges, allowing for exploitation of visual prior knowledge and training efficiency improvements.
  • Our proposed method exhibits competitive performance across different tasks in terms of both quantitative metrics and visual inspection, highlighting its effectiveness and generalization capability.
2 Related Works
2.1 Dense Prediction
Dense prediction refers to the type of problems whose outputs match the spatial dimensions of input data, often with the goal of assigning a value to each individual pixel. Typical examples include depth estimation, surface normal prediction, semantic segmentation, optical flow, etc. Common solutions can be categorized in discriminative and generative manner. Discriminative approaches focus on architecture design and introduce techniques like skip connections, dilated convolutions, multi-scale feature aggregation, and attention mechanisms [9, 44, 34] to better capture both local and global context for improved prediction accuracy. Generative methods, on the other hand, offer an alternative paradigm to tackle these problems through distribution modeling. Some works reuse or merge prior features from pretrained vision foundation models [4, 23, 62]. Others formulate these tasks as visual-condition guided generation process which refines the desired dense map iteratively [18, 32, 24].
2.2 Diffusion Bridge Models
Stochastic bridge models are a common tool in probability theory which describe the evolution of stochastic processes with given starting/ending points. When combined with diffusion models, they provide a new perspective to translate between different distributions and model conditional probabilities without the need for prior knowledge, therefore have been gaining increasing attention in the context of generative modeling in recent year [39, 38, 16]. They demonstrate the advantages of data-to-data generation paradigm over noise-to-data one on various tasks, including image-to-image translation [63, 37, 36], image restoration [59, 13], text-to-speech synthesis [10], etc. However, to the best of our knowledge, no prior work has explored leveraging diffusion bridge models on dense prediction tasks to facilitate the distribution transport between images and dense signal maps, despite their intuitive structural closeness. Our work aims to fill in this research gap.
3 Preliminary
3.1 Diffusion Process
A diffusion process =$\mathbb{Q}=\{\mathbf{z}_{t}\in\mathbb{R}^{d}:t\in[0,T]\}$ on space $\mathbb{R}^{d}$ follows the stochastic differential equation of form
$\mathrm{d}\mathbf{z}_{t}=f(\mathbf{z}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t}$ (1)
where $\mathbf{w}_{t}$ is a standard Wiener process, $f:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}$ and $g:[0,T]\rightarrow\mathbb{R}$ are the drift and diffusion coefficients that describe the evolution of this stochastic process. The transition kernel is specified as $q(\mathbf{z}_{t}|\mathbf{z}_{0})=\mathcal{N}(\alpha_{t}\mathbf{z}_{0},\sigma_{t}^{2}\mathbf{I})$ and the noise schedule comprising differentiable functions $\alpha_{t},\sigma_{t}$ depends on the design choice of drift and diffusion coefficients. It is also common to assess the diffusion process in terms of the signal-to-noise ratio (SNR), defined as $\text{SNR}_{t}=\alpha_{t}^{2}/\sigma_{t}^{2}$
3.2 Diffusion Bridge Process
Diffusion bridge process is a special type of diffusion process that guarantee to arrive at a prescribed final state $T$. One natural way of its construction is to derive the conditioned process from a reference diffusion process given terminal constraints. We can use Doob’s h-transform technique [42, 48] to modify a reference process $\mathbb{Q}$ defined in Eqn.(1) to ensure that the event $\mathbf{z}_{T}=\mathbf{x}$ happens. Therefore the conditioned process $\mathbb{Q}^{\mathbf{x}}(\cdot):=\mathbb{Q}\left(\cdot|\mathbf{z}_{T}=\mathbf{x}\right)$ follows the law of
$\mathrm{d}\mathbf{z}_{t}=\left[f(\mathbf{z}_{t},t)+g^{2}(t)h(\mathbf{z}_{t},t,\mathbf{x},T)\right]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t}$ (2)
where $h(\mathbf{z}_{t},t,\mathbf{x},T)=\nabla_{\mathbf{z}_{t}}\log q(\mathbf{z}_{T}=\mathbf{x}|\mathbf{z}_{t})$. Intuitively, the additional drift term $h(\mathbf{z}_{t},t,\mathbf{x},T)$ plays the role of steering $\mathbf{z}_{t}$ towards the specified target point $\mathbf{z}_{T}=\mathbf{x}$, and it is ensured that $\mathbb{Q}^{\mathbf{x}}$ invariably passes $\mathbf{x}$ at timestep $T$ with probability 1 for any initial state $\mathbf{z}_{0}$.
The transition kernel of $\mathbb{Q}^{\mathbf{x}}$, given by $q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$, can be subsequently derived through Kolmogorov equation:
$q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})$ $=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})$ (3)
$m_{t}$ $=\alpha_{t}\left(1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\right)$
$n_{t}$ $=\frac{\text{SNR}_{T}}{\text{SNR}_{t}}\frac{\alpha_{t}}{\alpha_{T}}$
$\sigma_{t}$ $=\sigma_{t}\sqrt{1-\frac{\text{SNR}_{T}}{\text{SNR}_{t}}},$
where $\text{SNR}_{t},\alpha_{t},\sigma_{t}^{2}$ are all ingredients defined in the reference process $\mathbb{Q}$.
4 Method
4.1 Problem Formulation
We formulate different dense prediction tasks like depth estimation, surface normal prediction, semantic segmentation etc., as image-conditioned generation problems and propose DPBridge to model the conditional distributions $p(\mathbf{y}|\mathbf{x})$, where $\mathbf{y}\in\mathbb{R}^{H\times W\times C}$ indicates the desired dense signal map (e.g. segmentation, depth, surface normal), and $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is an RGB image. In the following sections, we’ll first introduce the motivation of utilizing diffusion bridge models to handle dense prediction problems. Then, we’ll elaborate the details of our proposed DPBridge framework, which is constructed based on pretrained diffusion backbone through proper finetuning.
Figure 2: Overview of our training pipeline, which is conducted in the latent space based on diffusion bridge theory. To overcome the discrepancy with standard diffusion process and facilitate finetuning, we derive the maximum likelihood training objective of diffusion bridge process, propose a distribution aligned normalization strategy, and assist the training with an auxiliary image consistency loss.
4.2 Motivation: Bridge VS Diffusion
As introduced previously, the incompatibility that employ diffusion models to dense prediction task, which is essentially a image-conditioned generation problem, lies in its underlying iterative denoising process starting from Gaussian noise. In such cases, the distance between prior Gaussian distribution and data distribution remains large, requiring inevitably more sampling steps as well as ad-hoc conditioning mechanisms to converge to target distributions, at the risk of degraded performance and slow inference speed.
Bridge models, on the other hand, directly formulate generative processes as conditional probability paths. In this scenario, the boundary distribution is not restricted to Gaussian noise, but instead can be chosen as more informative priors, such as the condition data distribution. By establishing direct connections between different distributions, this formulation provides a more natural and efficient way to model probability transport compared with previous efforts. For example, it reduce the dependence on manually designed post-hoc guidance [11, 53, 12] or customized generation pipeline [41]. In addition, through implicitly integrates condition as part of the learned process, it provides a more flexible way to model intricacies between data and conditions without relying on explicit mechanisms like concatenation, attention [47, 61, 25], whose inductive bias may harm the overall performance when dealing with complex circumstances. Considering these advantages, we advocate that adopting diffusion bridge model will be a better fit to dense prediction tasks compared with conventional diffusion-based counterparts.
4.3 Dense Prediction Bridge
Given paired training data, a naive implementation is to construct the diffusion bridge model from scratch according to Sec.3.2. However, this approach is extremely resource-demanding, making it hard to implement with limited computation power. What’s more severe, it lacks the ability to harness rich scene understanding knowledge inherent in the vision foundation models, which have been proven significant in terms of boosting both perception quality and generalization performance in dense prediction tasks. Therefore, our DPBridge is build upon pretrained diffusion backbones through finetuning and we explore several strategies to handle the misalignment between original diffusion bridge process and standard diffusion process, aiming for more efficient knowledge transfer. Fig.2 gives an overview of the proposed finetuning pipeline, and we’ll explain the design of each component next.
4.4 Latent Space Implementation
Instead of directly learning in image pixel space, we follow common practices to perform diffusion steps in a compressed but perceptually aligned latent space, provided by an independently trained variational autoencoder (VAE). Specifically, both the target dense maps and RGB images are converted by the encoder to obtain $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$ and $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$, serving as the boundary states. In this way the diffusion bridge process is henceforth learned in the latent space. To retrieve original signal map we can use the decoder to decode the predicted latents $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$. In later part of this paper, with little abuse of notation, all derivations are conducted in the latent space.
4.5 Training Objective
4.5.1 Maximum Likelihood Loss
In previous diffusion bridge models, widely-used techniques for learning the reverse process that corresponds to Eqn.(2) is to follow the theory of [1], use a neural network to approximate the score function and apply score matching technique. We notice that analytical form of diffusion bridge’s reverse transition kernel is directly obtainable, therefore the usage of maximum likelihood training can also be applied, which is performed by optimizing the Evidence Lower Bound (ELBO). It yields a more stable loss function compared with score matching and better aligns with the pretrained backbone. Given the fixed terminal state $\mathbf{z}_{T}$, the expectation of data log-likelihood possesses the following ELBO formulation:
$\text{ELBO}=&\mathbb{E}_{p(\mathbf{z}_{0})}\Big{[}\mathbb{E}_{p(\mathbf{z}_{1}\mid\mathbf{z}_{0})}\big{[}\log p_{\theta}(\mathbf{z}_{0}\mid\mathbf{z}_{1},\mathbf{z}_{T})\big{]}-\\ \sum_{t=2}^{T}\mathbb{E}_{p(\mathbf{z}_{t}|\mathbf{z}_{0})}&\big{[}\mathbb{D}_{KL}\big{(}q(\mathbf{z}_{t-1}\mid\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\big{)}\big{]}\Big{]}\end{aligned}$
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ and $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ are the transition kernel of true reverse process and our modeled generative process and $\mathbb{D}_{KL}$ is the KL divergence. Utilize Bayes’ theorem and Markov chain property, we obtain:
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=\frac{q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})}{q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})}$ (5)
$=\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$
For simplicity, we define the following notations: $a_{t}=m_{t}/m_{t-1},b_{t}=n_{t}-a_{t}\cdot n_{t-1},\delta_{t}^{2}=\sigma_{t}^{2}-a_{t}^{2}\cdot\sigma_{t-1}^{2}$. The mean value term is:
$\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ $=k_{1}\mathbf{z}_{t}+k_{2}\mathbf{z}_{0}+k_{3}\mathbf{z}_{T}$ (6)
$k_{1}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}$
$k_{2}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot m_{t-1}$
$k_{3}$ $=\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}\cdot n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot a_{t}\cdot b_{t},$
and the variance is:
$\tilde{\sigma}_{t}^{2}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\cdot\delta_{t}^{2}$ (7)
Assume $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{T})$ is also a Gaussian distribution parameterized as $\mathcal{N}(\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))$ and let $\boldsymbol{\Sigma}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t))=\tilde{\sigma}_{t}^{2}$ for derivation simplicity, maximizing the ELBO is equivalent to minimizing:
$\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\frac{1}{2\tilde{\sigma}_{t}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{t}-\boldsymbol{\mu}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$ (8)
Similar to the strategy in DDPM [26], we use a network $\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)$ to parameterize noise instead of posterior mean and further simplified the loss as:
$\mathcal{L}_{elbo}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right],$ (9)
4.5.2 Image Consistency Loss
In addition to the common likelihood loss evaluated in the latent space, we also incorporate an image-level consistency loss, where we first convert the latent prediction back to image space, and then evaluate its consistency with the ground truth dense map using a simple MSE loss.
$\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y})\right\|^{2}$ (10)
The reason is that the compression of input data in the latent space will lead to a loss of high-frequency information, harming the model’s ability of preserving fine-grained spatial details. By introducing an auxiliary loss at image level, we ensure that the model is capable of retaining spatial information which is essential for precise per-pixel prediction. This dual loss strategy leverages the compactness of latent space for efficiency while still anchoring the training process to the fidelity of full-resolution dense map prediction, thus leading to performance improvement. Combined these two losses together, our final training objective becomes:
$\mathcal{L}=\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}$ (11)
where $\omega_{1},\omega_{2}$ are adjustable hyperparameters to balance the loss. An overview of DPBridge’s training pipeline shown in Algorithm.1.
Algorithm 1 Training
1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Loss weights $\omega_{1},\omega_{2}$
2:repeat
3:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
4:     $\mathbf{z}_{0}=\mathcal{E}(\mathbf{y})$, $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
5:     $t\sim\text{Uniform}(\{1,\dots,T\})$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
6:     $\mathbf{z}_{t}=n_{t}\mathbf{z}_{T}+m_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}$
7:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
8:     $\mathcal{L}_{elbo}=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t)\right\|^{2},$
9:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
10:     $\mathcal{L}_{ic}=\left\|\mathcal{D}(\hat{\mathbf{z}}_{0})-\mathbf{y}\right\|^{2}$
11:     Take gradient step on $\nabla_{\theta}\left(\omega_{1}\mathcal{L}_{elbo}+\omega_{2}\mathcal{L}_{ic}\right)$
12:until converged
Algorithm 2 Inference
1:Encoder $\mathcal{E}$; Decoder $\mathcal{D}$; Trained $\boldsymbol{\epsilon}_{\theta}(\cdot,\cdot,\cdot)$
2:$\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
3:for $t=T,\dots,1$ do
4:     $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\boldsymbol{\epsilon}=\mathbf{0}$
5:     $\mathbf{z}_{t}^{\prime}=\frac{1}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T})$
6:     $\hat{\mathbf{z}}_{0}=\frac{1}{m_{t}}(\mathbf{z}_{t}-n_{t}\mathbf{z}_{T}-\sigma_{t}\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t}^{\prime},\mathbf{z}_{T},t))$
7:     $\mathbf{z}_{t-1}=k_{1}\mathbf{z}_{t}+k_{2}\hat{\mathbf{z}}_{0}+k_{3}\mathbf{z}_{T}+\tilde{\sigma}_{t}\boldsymbol{\epsilon}$
8:end for
9:$\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
10:return $\mathbf{y}$
4.6 Distribution Alignment Normalization
A notable challenge during finetuning is the misalignment of intermediate sample’s marginal distribution $q(\mathbf{z}_{t})$ between DPBridge and diffusion backbone, due to the difference in the forward transition kernel. This misalignment will weaken the model’s inherent denoising ability, thus limiting the effectiveness of knowledge transfer. To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions. Specifically, we normalize the intermediate latents computed from Eqn.(3) to match the form of a standard diffusion process:
$\mathbf{z}_{t}^{\prime}=\frac{\mathbf{z}_{t}-m_{t}\mathbf{z}_{T}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}=\frac{n_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\mathbf{z}_{0}+\frac{\sigma_{t}}{\sqrt{n_{t}^{2}+\sigma_{t}^{2}}}\boldsymbol{\epsilon}_{t}$ (12)
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone.
4.7 Network Architecture
To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate and terminal state samples along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by [32], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
4.8 Accelerate Inference
Similar to the idea of DDIM [52], the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path. Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained. The accelerated inference process can be thus described as:
$\begin{split}\mathbf{z}_{t_{n}}&=n_{t_{n}}\mathbf{z}_{T}+m_{t_{n}}\hat{\mathbf{z}}_{0}+\\ &\sqrt{\sigma_{t_{n}}^{2}-g_{n}^{2}}\frac{\mathbf{z}_{t_{n+1}}-n_{t_{n+1}}\mathbf{z}_{T}-m_{t_{n+1}}\hat{\mathbf{z}}_{0}}{\sigma_{t_{n+1}}}+g_{n}\boldsymbol{\epsilon}\end{split}$ (13)
where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process. We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Eqn.(5), the process becomes a Markovian bridge. The basic inference process of DPBridge is summarized in Algorithm.2. If adopt acceleration, one can utilize Eqn.(13) during implementation.
5 Experiments
We evaluate the proposed DPBridge approach mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction. We’ll start by a brief introduction of experiment settings of these two tasks, followed by quantitative and qualitative analysis. Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim [45] and Virtual KITTI [8], which cover both indoor and outdoor scenes, for training. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 [51], KITTI [22], ScanNet [15], ETH3D [49] and DIODE [56]. For surface normal prediction, we additionally incorporate iBims [33] and Sintel [7] datasets in the evaluation pipeline. For fair comparison, we use the same validation split as in previous methods [32].
Evaluation Metrics. For depth estimation, we follow the affine-invariant evaluation protocol, where we first align the prediction results to the ground truth with least squares fitting, and then use Absolute Mean Relative Error (AbsRel) and $\delta_{1}$ accuracy to assess the performance. Denote the ground truth depth $\mathbf{d}$, predicted depth $\hat{\mathbf{d}}$, and the number of valid pixels $N$, the AbsRel is calculated as $\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{d}_{i}-\hat{\mathbf{d}}_{i}\|/\mathbf{d}_{i}$, and $\delta_{1}$ accuracy is measured using the ratio of pixels satisfying $\text{max}(\mathbf{d}_{i}/\hat{\mathbf{d}}_{i},\hat{\mathbf{d}}_{i}/\mathbf{d}_{i})<1.25$. For the metrics of surface normal task, we report the commonly used mean angular error between the ground-truth normal vectors and the prediction results, as well as the percentage of pixels with an angular error below 11.25 degrees.
5.2 Implementation Details
We implement DPBridge based on Stable Diffusion (SD) v1.5 [46] as backbone. During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 2 days to complete on a single Nvidia RTX 3090Ti GPU.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
[…]
Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms with best and second-best results marked by bold and underline. DM and BM correspond to diffusion models and bridge models respectively. All models are trained using SDv1.5 as backbone.
5.3 Quantitative and Qualitative Results
Tab.1 and Tab.2 summarize the quantitative metrics of DPBridge compared with diffusion-based and bridge-based counterparts, who are most closely related to our work. The reason we use SDv1.5 instead of more advanced SDv2.1 as backbone is that our current pipeline does not support v-prediction parameterization as in SDv2.1, and this incompatibility will undermine the model’s knowledge transfer ability. We’ll present more detail analysis in the ablation study. For a fair comparison, we re-implement all baseline methods using the same SDv1.5 backbone as we do, and report the results. We also conduct comparison with feed-forward and diffusion-based approaches based on SDv2.1 backbone, and the complete results is presented in the Appendix. DPBridge outperforms most methods in our evaluation setting and even be comparable with state-of-the-art that is trained with SDv2.1. Specifically, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge process indeed benefits the perception performance. In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
Figure 3: Qualitative comparison of monocular depth estimation methods on KITTI and ETH3D dataset. DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
Figure 4: Qualitative comparison between our (above) and Marigold trained on SDv1.5 (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
The qualitative results for depth estimation is conducted on KITTI and ETH3D datasets, as shown in Fig. 3. We can see that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range. In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows. We also present visualizations on in-the-wild images for surface normal prediction task. DPBridge is able to predict high-quality surface normals equipped with fine-grained textures. This strong generalization capability can be attribute to that we effectively harness extensive visual prior knowledge inherent in the pretrained diffusion backbone.
Method NYUv2 ScanNet iBims-1 Sintel
[…]
Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
DPBridge (Ours) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms with best and second-best results marked by bold and underline. DM and BM correspond to diffusion models and bridge models respectively. All models are trained using SDv1.5 as backbone.
Design Components NYUv2 KITTI
[…]
26.8 85.7 33.6 57.7
8.8 93.0 14.0 83.1
8.5 93.3 13.4 83.9
7.5 94.9 13.1 84.6
6.9 95.7 11.2 87.3
Table 3: The ablation study regarding the design components of DPBridge. FT stands for utilizing pretrained diffusion backbone during training. DAN denotes distribution aligned normalization technique. IC indicates image consistency loss.
Parameterization Backbone Steps NYUv2 KITTI
[…]
50 8.2 94.5 12.7 85.0
50 23.7 59.1 30.5 56.7
50 7.6 95.6 12.0 85.8
50 7.7 94.6 12.4 85.2
50 22.8 62.4 28.5 60.1
5 12.1 86.5 19.4 66.5
10 7.7 95.1 12.5 84.1
50 6.9 95.7 11.2 87.3
100 6.8 95.1 11.4 87.5
200 6.6 95.6 12.3 84.8
Table 4: The ablation study regarding the hyperparameter choice of diffusion bridge process. Options include different network parameterizations, the version of SD to use as backbone, and the number of denoising steps during inference.
5.4 Ablation Studies
To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets. The results are summarized in Tab.3 and Tab.4, with detailed discussions provided below.
5.4.1 Design Components
Finetuning VS Training From Scratch. To assess whether adapting pretrained diffusion backbones to our DPBridge improves perception performance, we compare the difference between finetuning and training from scratch paradigm. As shown in the first row of Tab.3, fine-tuning significantly outperforms training from scratch under identical experimental settings. This difference is due to two factors: first, foundation models trained on large, diverse datasets capture universal visual priors, aiding knowledge transfer to dense prediction tasks. Second, training from scratch requires more iterations for convergence. These observations suggest that building DPBridge on pretrained backbones is indeed advantageous.
Distribution Alignment Normalization. Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial. As shown in the second and fourth rows of Tab.3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
Image Consistency Loss. Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Tab.3 shows that while effective, its impact is less significant than that of distribution alignment normalization. This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data. However, when integrated with other design components, the overall performance is significantly improved.
5.4.2 Bridge Process Hyperparameters
Parameterization and Backbone Selection. Apart from the noise prediction scheme used in our main experiments, we also consider how the choice of parameterization combined with the backbone selection influences performance. Results in Tab.4 reveal two key findings: (1) v-parameterization performs significantly worse compared with other two options, highlighting the need for further research on deriving objectives specific to the diffusion bridge process, and (2) ”noise prediction + SDv1.5” achieves the best results, indicating that the parameterization during fine-tuning should align with that of the backbone to maximize knowledge transfer efficiency.
Number of Sampling Steps. We evaluate performance with varying numbers of sampling steps during inference. As shown in the lower part of Tab.4, increasing sampling steps initially improves perception performance rapidly before leveling off. Notably, compared to the results in Tab.1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
6 Conclusion
In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks. We establish a direct mapping between input RGB images and their corresponding dense maps in the latent space based on fully-tractable diffusion bridge process. In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones. In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance. Future research directions include deriving v-prediction parameterization of diffusion bridge process, and extending to other dense prediction tasks.
Supplementary Material
In the supplementary material, we’ll provide contents that are not included in the main paper due to page limitations. We first fill in the derivation details of the loss function. Afterwards we will provide the additional quantitative comparisons with more baseline methods. Finally, we will present qualitative examples that apply DPBridge on more dense prediction tasks.
Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Sec.4.5.1 . Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Eqn.(5) , whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Eqn.(3)
$\begin{split}q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T},\sigma_{t}^{2}\mathbf{I})\\ \mathbf{z}_{t}&=m_{t}\mathbf{z}_{0}+n_{t}\mathbf{z}_{T}+\sigma_{t}\boldsymbol{\epsilon}_{t}\end{split}$ (14)
$\begin{split}q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})&=\mathcal{N}(m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T},\sigma_{t-1}^{2}\mathbf{I})\\ \mathbf{z}_{t-1}&=m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}+\sigma_{t-1}\boldsymbol{\epsilon}_{t-1}\end{split}$ (15)
To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Eqn.(14) with derivation from Eqn.(15), obtain
$\mathbf{z}_{t}$ $=a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T}+\delta_{t}\boldsymbol{\epsilon}$ (16)
$a_{t}$ $=\frac{m_{t}}{m_{t-1}}$
$b_{t}$ $=n_{t}-\frac{m_{t}}{m_{t-1}}n_{t-1}=n_{t}-a_{t}n_{t-1}$
$\delta_{t}^{2}$ $=\sigma_{t}^{2}-\frac{m_{t}}{m_{t-1}}n_{t-1}^{2}\sigma_{t-1}^{2}=\sigma_{t}^{2}-a_{t}^{2}\sigma_{t-1}^{2}$
Thus, we can get $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})=\mathcal{N}(a_{t}\mathbf{z}_{t-1}+b_{t}\mathbf{z}_{T},\delta_{t}^{2}\mathbf{I})$, which is also a Gaussian distribution.
Putting all these terms back to Eqn.(5) , we’ll have:
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\sim\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t}(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T}),\tilde{\sigma}_{t}^{2}\mathbf{I})$ (17)
$=$ $\frac{q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})}{q(\mathbf{z}_{t}|\mathbf{z}_{0},\mathbf{z}_{T})}$
$\propto$ $\exp\bigg{[}-\frac{1}{2}\bigg{(}\frac{(\mathbf{z}_{t}-a_{t}\mathbf{z}_{t-1}-b_{t}\mathbf{z}_{T})^{2}}{\delta_{t}^{2}}+\frac{(\mathbf{z}_{t-1}-m_{t-1}\mathbf{z}_{0}-n_{t-1}\mathbf{z}_{T}^{2}}{\sigma_{t-1}^{2}}-\frac{(\mathbf{z}_{t}-m_{t}\mathbf{z}_{0}-n_{t}\mathbf{z}_{T}^{2}}{\sigma_{t}^{2}}\bigg{)}\bigg{]}$
$=$ $\exp\bigg{[}-\frac{1}{2}\bigg{(}\bigg{(}\frac{a_{t}^{2}}{\delta_{t}^{2}}+\frac{1}{\sigma_{t-1}^{2}}\bigg{)}\mathbf{z}_{t-1}^{2}-2\bigg{(}\frac{a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}}{\delta_{t}^{2}}+\frac{m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}}{\sigma_{t-1}^{2}}\bigg{)}\mathbf{z}_{t-1}+C(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})\bigg{)}\bigg{]}.$
where $C(\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ indicates the constant term that is independent of $\mathbf{z}_{t-1}$. Since the Gaussian distribution probability density $\mathbf{z}_{t-1}\sim\mathcal{N}(\tilde{\boldsymbol{\mu}}_{t},\tilde{\sigma}_{t}^{2}\mathbf{I})$ can be expressed in the following form:
$\exp\left(-\frac{(\mathbf{z}_{t-1}-\tilde{\boldsymbol{\mu}}_{t})^{2}}{2\tilde{\sigma}_{t}^{2}}\right)=\exp\left[-\frac{1}{2}\left(\frac{1}{\tilde{\sigma}_{t}^{2}}\mathbf{z}_{t-1}^{2}-\frac{2\tilde{\boldsymbol{\mu}}_{t}}{\tilde{\sigma}_{t}^{2}}\mathbf{z}_{t-1}+\frac{\tilde{\boldsymbol{\mu}}_{t}^{2}}{\tilde{\sigma}_{t}^{2}}\right)\right].$ (18)
We can thus derive:
$\frac{1}{\tilde{\sigma}_{t}^{2}}$ $=\frac{a_{t}^{2}}{\delta_{t}^{2}}+\frac{1}{\sigma_{t-1}^{2}}$
$\Rightarrow\tilde{\sigma}_{t}^{2}$ $=\frac{\sigma_{t-1}^{2}(\sigma_{t}^{2}-a_{t}^{2}\sigma_{t-1}^{2})}{\sigma_{t}^{2}}=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\delta_{t}^{2}$ (19)
$\frac{\tilde{\boldsymbol{\mu}}_{t}}{\tilde{\sigma}_{t}^{2}}$ $=\frac{a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}}{\delta_{t}^{2}}+\frac{m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}}{\sigma_{t-1}^{2}}$
$\Rightarrow\tilde{\boldsymbol{\mu}}_{t}$ $=\bigg{[}\frac{\sigma_{t-1}^{2}\big{(}a_{t}\mathbf{z}_{t}-a_{t}b_{t}\mathbf{z}_{T}\big{)}+\delta_{t}^{2}\big{(}m_{t-1}\mathbf{z}_{0}+n_{t-1}\mathbf{z}_{T}\big{)}}{\delta_{t}^{2}\sigma_{t-1}^{2}}\bigg{]}\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\delta_{t}^{2}$
$=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}\mathbf{z}_{t}+\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}m_{t-1}\mathbf{z}_{0}+\bigg{(}\frac{\delta_{t}^{2}}{\sigma_{t}^{2}}n_{t-1}-\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}a_{t}b_{t}\bigg{)}\mathbf{z}_{T}$ (20)
whose results are equivalent to Eqn.(6) and Eqn.(7).
Accordingly, maximizing the Evidence Lower Bound (ELBO) is equivalent to minimizing the KL divergence between ground truth and predicted posterior:
$KL\left(q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})\,||\,p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},\mathbf{z}_{T})\right)$ (21)
$=$ $\mathbb{E}_{q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})}\left[\log\frac{\frac{1}{\sqrt{2\pi}\tilde{\sigma}_{t-1}}e^{-\frac{(\mathbf{z}_{t-1}-\tilde{\boldsymbol{\mu}}_{t-1})^{2}}{2\tilde{\sigma}_{t-1}^{2}}}}{\frac{1}{\sqrt{2\pi}\tilde{\sigma}_{\theta,t-1}}e^{-\frac{(\mathbf{z}_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1})^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}}}\right]$
$=$ $\mathbb{E}_{q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T})}\left[\log\tilde{\sigma}_{\theta,t-1}-\log\tilde{\sigma}_{t-1}-\frac{(x_{t-1}-\tilde{\boldsymbol{\mu}}_{t-1})^{2}}{2\tilde{\sigma}_{t-1}^{2}}+\frac{(x_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1})^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}\right]$
$=$ $\log\frac{\tilde{\sigma}_{\theta,t-1}}{\tilde{\sigma}_{t-1}}-\frac{1}{2}+\frac{\tilde{\sigma}_{t-1}^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}+\frac{(\tilde{\boldsymbol{\mu}}_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1})^{2}}{2\tilde{\sigma}_{\theta,t-1}^{2}}$
If we assume that only the mean term is learnable during training, and $\tilde{\sigma}_{\theta,t-1}=\tilde{\sigma}_{t-1}$, we can ignore all the unlearnable constants and arrive at the final training objective:
$\mathcal{L}$ $=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\frac{1}{2\tilde{\sigma}_{t}^{2}}\left\|\tilde{\boldsymbol{\mu}}_{t-1}-\tilde{\boldsymbol{\mu}}_{\theta,t-1}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$ (22)
$=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{z}_{T}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{z}_{t},\mathbf{z}_{T},t)\right\|^{2}\right]$
which is the same as in Eqn.(8) and Eqn.(9) . This concludes our derivation.
Appendix B Additional Quantitative Results
Next, we provide additional quantitative results on depth estimation (Tab.5) and surface normal prediction (Tab.6) tasks. Compared with Tab.1 and Tab.2 presented in the main paper, we conduct comparisons with more baseline methods, including feed-forward approaches and diffusion-based methods that is trained on more advanced Stable Diffusion (SD) v2.1 backbone. In depth estimation, DPBridge consistently matches state-of-the-art methods, while demonstrating robust generalization across datasets. For example, it achieves 6.9 AbsRel, 95.7 % $\delta_{1}$ on NYUv2, and 7.5 AbsRel, 94.7% $\delta_{1}$ on ScanNet. Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 18.1 and $11.25^{\circ}$ accuracy of $54.2\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 24.3, $11.25^{\circ}:51.2\%$) and Sintel (mean: 41.2, $11.25^{\circ}:19.4\%$) and. These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
[…]
DiverseDepth [57] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1
MiDaS [43] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5
LeReS [58] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6
Omnidata [19] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2
HDN [60] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0
DPT [44] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8
Marigold* [32] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2
DepthFM* [24] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0
Geowizard* [20] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1
E2E-FT* [21] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6
Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
Table 5: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
Method Model Type NYUv2 ScanNet iBims-1 Sintel
[…]
EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5
Omnidata v1 [19] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4
Omnidata v2 [31] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7
DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5
Marigold* [32] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9
Geowizard* [20] 19.1 49.5 17.3 53.7 19.5 61.6 40.4 13.2
E2E-FT* [21] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3
Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
DPBridge 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
Table 6: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
Furthermore, we also evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise. The evaluation is conducted under both depth estimation and surface normal prediction tasks. We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1. Tab. 7 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise. Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity gaussian noise cause the most significant performance decline. In addition, the noise intensity has more impact on surface normal prediction than depth estimation, showcasing the inherent characteristics between different tasks. These results emphasize the model’s robustness under moderate noise but still expose its vulnerability to extreme or specific types of noise.
Noise Type Noise Level Depth Estimation Surface Normal Prediction
NYUv2 KITTI NYUv2 iBims-1
[…]
Original Version - 6.9 95.7 11.2 87.3 18.1 54.2 24.3 51.2
Gaussian $\mathcal{N}(0,0.05)$ 7.2 94.4 11.8 86.0 18.6 52.2 25.6 45.1
$\mathcal{N}(0,0.1)$ 8.4 92.7 12.1 85.3 20.3 49.4 27.4 40.4
$\mathcal{N}(0,0.2)$ 10.3 90.0 15.7 78.6 24.6 43.7 37.0 25.4
$\mathcal{N}(0,0.5)$ 15.9 77.2 22.5 67.1 41.3 22.5 50.0 13.2
Uniform $\mathcal{U}(-0.05,0.05)$ 7.0 95.0 11.4 86.5 18.3 53.6 24.5 50.7
$\mathcal{U}(-0.1,0.1)$ 7.2 94.7 11.5 85.4 18.7 53.0 24.8 49.4
$\mathcal{U}(-0.2,0.2)$ 7.6 94.0 12.0 85.0 19.5 50.5 26.0 46.1
$\mathcal{U}(-0.5,0.5)$ 9.3 91.1 16.3 76.1 22.4 42.2 30.2 31.6
Poisson Poisson($\lambda=0.05$) 10.2 90.2 13.1 82.5 21.3 45.6 28.1 38.9
Poisson($\lambda=0.1$) 11.8 88.9 15.4 79.3 23.2 40.3 34.9 24.8
Salt & Pepper Probability=5% 12.0 85.4 13.3 81.7 26.3 28.8 31.8 29.6
Probability=10% 14.3 82.2 16.0 76.2 28.6 26.3 35.3 21.4
Table 7: Robustness evaluation on depth estimation and surface normal prediction tasks with different types and different intensities of noise. During all experiments, the input images are normalized between -1 to 1.
Appendix C Additional Qualitative Results
To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems. For semantic segmentation, we convert discrete labels into color maps and use our bridge model to predict colors in RGB space instead of categorical labels. For optical flow, the bridge process is defined between the flow image and the first image in the input RGB pair. To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates. The following figures (Fig.5, Fig.6, Fig.7, Fig.8) illustrate that DPBridge achieves visually satisfactory results across different tasks. Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
Figure 5: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [14] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [35] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 7: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [7] dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 8: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [55] dataset.